MechForce

Version 3.80

14 October 1993

edited by Andreas Smigielski, Wiedweg 5, 38120 Braunschweig, Germany

by Ralph H. Reed P.O. Box 1497, Eglin AFB, FL 32542, USA

1 Copyright and Disclamier notice

MECHFORCE \begin{matrix}c \\ ◯ \end{matrix}1989 By Ralph H. Reed

This software is very similar to the BattleTech system which is copyright 1985-1989 \begin{matrix}c \\ ◯ \end{matrix} FASA Corporation. However it not exactly the same. Some ideas have been used without permission of FASA Corporation. Some areas and ideas have been changed to protect the rights of FASA Corporation and others. The Mechforce program, its code, ideas, pictures and documentation within are Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989 by Ralph H. Reed.

MechForce is a utility system which allows people to test "mechs" in a "game arena" for the purpose of seeing how they would do against other mechs. Previous version of BattleFort and BattleForce are no longer supported.

2 What is MechForce?

It is a game that simulates combat between two or more giant, robot-like machines. Simple words can’t begin to give you the feel of piloting a 30 - 40 foot tall, fire breathing, earth shaking colossus that obeys your every whim. To enjoy the game you must put yourself in the cockpit of a Combat Mech. Live the experience!

You are the elite of the elite - a Mech Warrior. Your Combat Mech thunders under you as you pace across the land, searching for a reported enemy mech. You view the land around you, searching for the enemy. To your left the plain climbs gradually from low hills up to high mountains. To your right is a lake, the calm waters shading from light to cobalt blue as the water deepens in the center, with patches of treacherous swamp at the edges. Ahead of you the hills meet a mixture of light and heavy woods. You are debating whether to climb the hills or fight your way through the trees when the enemy mech steps over the crest of the hill.

Your heart leaps as you recognize a heavy Archer, just as he fires one of his massive 20 racks of Long Range Missiles. Your Marauder is staggered by the multiple impacts as you frantically pit your piloting skill against 75 tons of overbalanced mech. You swivel your torso and call up your weapons display. The reassuring green lights indicate that all weapons are ready to fire. You select your Auto Cannon 5 and one of your Particle Projection Cannons. A moment later your combat display shows you your field of fire and weapon range. The targeting pipper is blinking on the Archer as you fire your weapons up the mountain.

The crash of your cannon tears chunks of armor from the Archer, but the scintillating beam of the Plasma Cannon slashes to the left of the Archer. "Too quick" you curse under your breath. The temperature in your cockpit jumps as the heat from firing the Plasma Cannon soaks into your mech. No problem so far, the heat sinks can handle it. You drive your mech into a pounding run. You’ve got to get close, his LRMs have a minimum effective range, if you can just get inside it...

The Archer is shrouded in smoke as another salvo of missiles blast their way toward you. Missiles clang off your armor, tearing chunks away. Your cockpit rings as one hits the head of your mech. You fight for consciousness and keep going, his racks are empty now, he has to reload.

Motion, behind you. Another Combat Mech breaks from the forest and races across the plains toward you. You quickly identify it as a light weight Stinger, ordinarily no threat to a Marauder, but it AND an Archer? What to do? Run? Fight? Fight which one?

There’s no real choice, if you turn your back on the Archer he’ll nail you. Ignore the Stinger. You quickly check your status. Your heat is dropping as the heat sinks suck away the excess heat. Your armor indicators are still all green although those missiles have chewed away about half of you left leg armor. Your Auto Cannon is still recycling, but your Plasma Cannons are ready. Time to Rock an’ Roll.

Weapons select. You arm both of your Plasma Cannons. The targeting computer comes up and you select the Archer again. Your mech is rocked by a laser blast in its vulnerable back, but you ignore it, triggering your Plasma Cannons. The warbling shreik of your Plasma Cannons echos across the plains. Blazing energy lances from each arm of your mech to impale your enemy. The first strikes the Archer dead center where his armor is the strongest, but the second beam scores a lucky hit on his vulnerable head. The Archer is staggered as armor vaporizes from his chest and head. You watch the Mech fight for balance and fail, tumbling backward off the mountain.

One down. Another laser digs into your back, as red lights flash on your status display. Your back armor is breached, the laser got through to your Auto Cannon and its out of the action. You begin to turn your ponderous steed. As you turn to face the charging Stinger it suddenly jumps into the air on jets of flame.

"Damn!!!" you yell as you frantically bring up your weapons. The Auto Cannon in your Left Torso is disabled, your Plasma Cannons are still recycling. You’re still hot, but there’s no choice as you arm the two medium lasers in your chest. You try to target the fast moving Stinger, but before you can fire 20 tons of armor and fire power land on top of you.

Your mech shutters under the impact. You scramble to control it, but you’re falling. "The next time I’m in the Factory I’m putting jump jets in this thing" you vow. The ground shakes as 75 tons of Mech falls flat on its back. Lights are flashing all over. You bring up your status display. The Stinger landed on your right shoulder, its out along with all the weapons in your right arm. The fall did more damage to your weak left leg. You’ll survive, now where is that little Stinger. You look around. The Stinger is racing for the forest.

"Oh no you don’t" you mutter as you wait for your remaining Plasma Cannon to recycle. You target the Plasma Cannon at the racing Stinger’s back. The warbling blast echoes across the plain as your man-made lightning bolt hits the Stinger dead center.

The Stinger stumbles, starts to fall... a blinding flash lights the sky as its fusion reactor explodes. Metal peppers your mech as the concussion threatens to knock you unconscious. Your temperature gauges peg as the mech shuts down from too much heat. With time your mech will recover. Lets just hope the Archer has had enough.

This is MechForce. It can be played with one person against the computer or up to 32 people fighting in teams. You can train in the simulator, fight in the arena, or try to complete combat missions.

3 MechForce Configuration

The MechForce program use a config file located in the current directory. The config file named "MechForce.cfg" will preset several variables for the game and all of the paths for the many files that the program use. If the config file can not be found the program will default to directories in the current directory, it will also default to save turned off.

You can change the included config file with the included "ConfigEd" program. Make sure the paths are valid paths and correct. THIS IS VERY IMPORTANT. If the program can not find the files it need they will not work.

The MechForce program will look for images files in a subdirectory called "Images" in the current directory. It will look for sound files in "Sound", map files in "Maps", saved games in "Games", unit files in "Units", warrior and technician files in "Warriors" and mech files in the "Mechs" directory. In addition, it expects mechs to be subdivided into four classes of mechs, "Light" 20-35 tons, "Medium" 40-55 tons, "Heavy" 60-75 tons, and "Assault" 80-100 tons. If you create new mechs and put them in the wrong subdirectory some parts of the programs will have trouble finding your new mechs.

4 A word or two about memory

MechForce is a huge program. If your system has only 512K of memory you will be limited in the number of mechs you can load. Computer mechs take a lot more memory than player mechs. You will probably be able to load 24 to 32 player mechs, but if you are playing against the computer 8 player and 8 computer will probably be the limit. If your load too many mechs there will not be enough memory left to load the sound files.

5 MechForce Documentation

The first thing you will see when the game starts is a requester asking you to select the Type of Game. The first option allows you to Continue a Saved Game. The second option is a Simulation mode and is used for training. This mode allows all combinations of teams and mech types, but damage and experience are not saved at the end of the game. At the start of the game you will be shown a list of available orders and you can pick the set of orders to follow. If you hit CANCEL you will be training for the Roving Patrol mode. Arena Challenge is a One on One combat playing against another person or the computer. If the computer is the opponent it will pick its own mech. Arena mode will also let you Bet on the outcome of the combat. One of Major Houses may offer to Sponsor you in the Arena. This may offer a little protection from other friendly houses, but it will make you a target of enemy houses. Not having a sponsor won’t keep some houses from picking on you. This takes the form of Sabotage to your mech. The computer will check and see if your pilot can detect any sabotage. If he does he will be given the option to try to repair the sabotage. Sabotage may make a critical part fail during combat. Trying to repair the sabotage may destroy it completely. The chance of sabotage gets higher as you win more Arena Combats. Roving Patrol is a general purpose combat option. There is no goal other than to kill the enemy and survive. It is real combat and you do earn a few Combat Points. Combat Orders is like the Simulation mode, all combinations of teams and mech types are allowed, the difference is that damage and experience are saved at the end of the game. The main difference is that the computer will pick the orders randomly for your team to follow.

Once you have selected the type of game you want to play you will get a Team Selector. The Team Selector window contains four rows of selector boxes. The center box contains the Team number. If you are playing Arena or Patrol mode, only the first two rows will be active. The Left and Right boxes in each row represent the two sides of a combat. To put a Team on a side just click in the box. The word "Team" and the team number will appear in the box. To change sides, click on the other side. To remove a team, click on the same side a second time. Once a team is placed on a side the central box becomes the selector for the teams control device. An "M" will appear indicating that the default control device is the Mouse. By clicking on the center box you can also select "J" for Joystick, "K" for Keyboard and "C" for Computer controlled. If computer control is selected, ALL teams on that side will be computer controlled.

The next thing you need for combat is a place to fight. If you are in any of the combat modes the computer will randomly select one of seven special maps. In Training or Orders mode the map will be generated based on the type of map called for in the orders. If "Create" is called for in the orders you will get the Random Map Control Panel See Chapter 10 [The Random Map Control Panel], page 9.

The next thing you need are mechs to fight. Two windows will open. The first one shows the Team you will be working with in the central box on the bottom. The second window is a file requester for loading mechs. There are three flavors of mechs. Generic "Mechs" are the type you create in the HeadQuarters program with "Create a Spare Mech". The pilot for these mechs will be randomly created by the computer. "Warrior" files are created in the HeadQuarters program and should contain a mech and a pilot. "Unit" files are collections of warriors, mechs and technicians all in one common file.

Only Simulation mode will allow a player to fight with a plain mech, all other modes require you to use a Warrior file or a warrior from a Unit file. All members of a team MUST be of the same mech type. Warriors are marked when they are loaded and can not be loaded twice. If you save a game in progress those warriors can not be used in another game until the first game is completed.

To load a file, click on the file you want and then click on "Load". If the file is a plain mech you will get a requester asking you to name the pilot. To load warriors from a Unit, Load the Unit and then select the warrior from the list of warriors in the unit. The mechs and pilots you selected will appear in the Team window. When your team is complete, click on the "Done" gadget in the file requester window and it will disappear. At this point you can use the small selector boxes to the right of each name to customize the control device for each player.

At any time your can click on a mech name with the RIGHT mouse button and get a quick preview/status display of that mech. This can help you avoid loading a damaged mech. You can not change the device for a computer mech. To change teams, click the "Team" gadget. The "Load Mechs" gadget will bring the file requester back. When you are done with all teams, click the "Done" gadget in the team window. You will also notice a small gadget in the upper right corner with an "R" in it. If you make a mistake loading mechs you can click on the Reset gadget and it will clear ALL mechs. Then you can start over in your selection.

All of the windows leading up to and including the Team window have an ESCAPE option. If you hit the ESCAPE key you will be asked if you want to quit the game. If you say YES it will exit back to WorkBench.

From this point on everything is controlled from a control menu at the bottom of the screen. These menus consist of a prompt and two or more options. The selected option will be highlighted. Control depends on the device you are using. The Keyboard uses the Left and Right arrows to change and the Space Bar to select. The Joystick uses Left and Right stick command to change and the Fire Button to select. The Mouse changes by moving the mouse over the desired option and selects with the Left Button. The mouse will only work when it is positioned in the menu line.

The menus are pretty much self explanitory. There are two main menus associated with two timers. One for Movement and one for Weapons. Everything is based on TIME. Since there are TWO timers you can move and fire weapons at the same time. The Wait options are especially powerful if used properly because you can wait in weapons and still move.

6 Movement Menu

• Walk: Walk forward or backward one hex. Takes base time + heat penalty + damage penalty. The base movement time and penalty times can be found in the status display. Walking generates 1 pt of heat/second. You can Crawl (Walk) while you are down at double the normal time.

• Run: Run forward. Takes 2/3 the time of Walk and generates 2 pts of heat/second.

• Jump: Use jump jets to jump in any direction. When you select Jump you will be asked for the direction. "E" is to the right and "W" is to the left. Once you have started jumping you can only continue jumping in the same direction or one hex face to either side. If you are already jumping you will get a different menu that allows you to pick the next direction, turn, stop or check options. You can jump approximately the same number of hexes you can walk in 100 seconds. This is a jump time of about 67 seconds. This may be less if you have fewer jump ports than the maximum allowed.

Every time you damage a jump port you loose one hex of range. Remember that you jump slower if you are hot so you may not get your max range. If you are still jumping when you run out of jump time you will CRASH!!! Once you land you can not jump again for 100 seconds from the time you STARTED jumping. The main menu will say "Jump (OK)" if you are ready to jump. It will say "Jump (No)" if your mech can not jump. It will say "Jump (xx)" with the number of seconds until your jump jets are ready again, if you are recycling.

• Turn: Change direction 60 degrees (one hex side). Time is based on whether you are standing, walking or running. Turning while jumping is handled in the jump menu and takes no time. You can turn while you are down at double the normal time.

• Stand: Stand up after your mech has fallen down. Standing takes the same time as moving one normal hex.

• Wait: Wait is used to control the timing of events in the Move timer. You can wait On Time, wait until you Call Move, wait until all of your weapons are recycled or the next weapon is recycled or your arms&legs are recycled, wait until your heat drops to a certain point (color), wait on your weapon selection, or wait on your enemies move. If no enemies are visible you will be given the option to wait on the next mech to become visible. You can chose to wait on the next Visible mech with the ’Visible’ option. Beware, if all mechs are visible this acts the same as a Wait On Call.

• Call Attack: This option will call the attack menu if it is waiting. You can also ’Call Attack’ from the Options menu and interrupt any attack WAIT command. ’Call Attack’ sets a flag, but leaves you in the move menu.

• Options: This option brings up another menu of special options.

• Status: Status will bring up your status display window.

• Identify: The Identify option will allow you to click individual mech images to identify their type and weight. You have to option for a full Status Display.

• Break Attack: Break Attack will interrupt an attack in progress.

• Break Move: The Break Move option will interrupt a move in progress.

• Special: The Special option calls another menu with five more options.

• Drop Ammo: The Drop Ammo option will ask you if you want to drop each ammo rack in your mech.

• Eject: Is used to leave a badly damaged mech. There is a skill roll for Ejecting from a mech. If you fail your pilot takes damage. You eject into the hex in front of your mech. DON’T eject into a forest fire. You will also have to make a Survival roll to keep from being captured.

• Surrender: This means that you stay with your mech, but you have quit fighting. If your side loses captured mechs and pilots belong to the winning side.

• Self-Destruct: The Self-Destruct option makes the thermonuclear engine of your mech explode. This option can only be used by Player characters that have advanced beyond the "Green" rank. You can not Surrender or Eject of you are Unconscious or your mech is Shutdown due to heat.

• Orders: The Orders option will give you you a brief review of your Orders.

Options do not take any time.

7 Weapon Menu

• Weapon: Brings up the weapon select display. This display shows all weapons, including arms and legs, where they are located ( LT=Left Torso, etc), and their condition. If the status indicator is green they are ready to use. Red indicates damaged or out of ammo. Yellow means they are recycling and is a bar graph of how long until ready to use again.

• Wait: Use this to control the timing of events in the Weapons timer. You can wait On Time, wait On Call, wait on your current move, wait until all your weapons are recycled or your first weapon to recycle or your arms&legs are recycled, wait until your heat drops to a certain color, or wait On Enemy move.

• Call Move: This option will call the movement menu if it is waiting. You can also ’Call Move’ from the Options menu and interrupt any movement WAIT command.

• Options: This is the same as the Move Options except the time comes off the weapons timer.

The movement and weapons timers are separate so you can move and fire weapons simultaneously. The wait conditions allow rapid movement without bothersome weapons messages, or standing still and firing, without bothersome move messages. For even more accurate timing you can use the Break feature and then Break Move or Break Attack from Options menu item. The Options items are also available as pulldown menus at the top of the screen. There is a small experience penalty for using the Break Attack and Break Move options from the pulldown menus. You can still shoot at Surrendered, Ejected and Dead mech. You can not shoot at Escaped or Destroyed mechs.

8 Pulldown Menus

• Game Menu: This menu has six items on it.

• About: Tells you a little about the Author, me.

• Save: Gives you two submenus, "Game" and "Map". Save Game will save the game at the end of the current second. Save Map will save the game map only to a file that can be used in another game.

• Sound: Toggles the sound on and off. If you didn’t have enough memory to load the sound files this will start as "Silent".

• Setup: Brings up a window that lets you change the speed of missiles and the delay for messages. It will also display the name of the first player and his Input Device. By Clicking on the Name you can flip through each player. By Clicking on the Device you can flip through the devices. This allows you to change player Input Devices during the game.

• Color: Changes the color of Mechs from Red to White. This if for some people that have trouble telling the difference between red and green.

• Quit: Allows you to quit a game in progress. For most real combats this option is disabled.

• Team Menus: Each player team has a pulldown menu with all of the members of that team as items. If the team is computer controlled you must use the "Identify" option to find out about enemy mechs. If the computer mech is not visible on the screen you will not be able to identify its status display. There is a small experience penalty for using the "Break Attack" and "Break Move" items from the pulldown menus.

9 Skill Rolls

While moving or in combat you will occassionally be asked to make a skill roll. This means that something has made your mech off balance (slipping in water, heavy weapons hit) and you as a pilot must roll to see if your skill is sufficient to keep from falling down. Falling in heavy mech is not trivial. The mech will take damage equal to its weight divided by 10. If you roll down a mountain this is multiplied by the number of levels you fall. Landing in water cuts your damage in half. All damage is applied in groups of 5 points all over your body. Landing on your head can ruin your day. Crashing is the same as falling only from three levels up (triple damage). When you are asked to make a skill roll there will be an option to "Modify" your skill roll. If you select this option you can trade experience points for an improved skill roll. This is expensive, but in critical situations it could save your life. If you use it too much you will not have any experience to improve your character. The computer can also modify its rolls and you may see a "modified+2" occasionally.

Firing weapons requires a skill roll vs your Gunnery skill to hit your target. You can modify this roll also. If you miss a shot there is a chance that you might hit someone else in an adjacent hex. There are also skill rolls for being knocked unconscious, shutting down due to heat and ammo explosions due to heat.

10 The Random Map Control Panel

The Random Map Control Panel is a little complicated. It contains six slider gadgets. The top three control the percentage of Forest, Mountains and Water in your map. The bottom gadget controls the density of the top three. A cluster is one group of woods or mountains or water. It can be one hex or many hexes. The more clusters you have the more details you will have in your map. The Sand and Snow requesters control the amount of sand and snow in you map. Sand and snow are added by replacing whatever plain hexes are left after the forest, mountains and water hexes are added. The gadget controls the percentage of plain hexes that will be replaced. Sand replaces hexes in the bottom half of the map and snow replaces hexes in the top half. Once you have everything set the way you want it, click the "Create" gadget and your world will be drawn. A small window will then appear asking you to comfirm your world, "OK", or try again, "ReDo".

11 Terrain Types and Features

• Plain: This is basic ground level, open plain, and is what all movement times are based on. The base movement time for a Plain hex is given in the status display.

• Sand: It takes three times as long to move thru sand and it adds 2 pts of heat per second.

• Snow: It takes twice as long to move thru snow, and there is a chance that you will slip on the ice. Snow will remove 2 pts of heat per second.

• Light Forest: It takes twice as long to move through light forest as across plains, and makes you one point harder to hit.

• Heavy Forest: It takes three times as long to move through heavy forest and you are two points harder to hit.

• Jungle: You can not move thru this hex. It is possible to jump over it, but if you land in a Jungle hex the only way out is to jump. Jungle hexes can be burned.

• Swamp: It takes twice as long to move, you are one point easier to hit, and there is a good chance of slipping and falling.

• Shallow Water: This type hex takes twice as long to move through and makes it easier for you to be hit and harder for you to hit others, but helps you to cool down. The water will remove 3 points of heat per second in addition to your heat sinks. There is a chance you will slip and fall.

• Medium Water: This type water is up to the mechs waist. Weapons below the waist can not be fired. You get 6 points of cooling per second, but you are easier to hit and it is harder for you to aim. You can’t kick in medium water. Move at 1/3 normal rate.

• Deep Water: This type water provides 9 points of cooling/second. Weapons below the neck can not be fired and you are very hard to hit. You can’t kick or punch in deep water. Move at 1/4 normal rate.

• Bottomless Water: This type water is beyond your mechs capability. It will crush your mech, destroying it and killing you. The only way past this type of hex is to jump over.

• Low Hills: Hills provide cover. You can’t shoot over them. Low hills make it one point easier to hit mechs below you. Move at 1/3 normal rate uphill or half normal rate on the same level.

• Medium Hills: You get two extra points to hit mechs below you due to height. Move at 2 + 1/level going uphill.

• High Hills: You get three extra points to hit mechs below you due to height. Hills have a tactical advantage, but beware of skill rolls. If you fall you can roll all the down, doing multiple damage.

• Barrier Peak: This hex type is impassible. You can not jump over it. It is used for obstacles. All Arena maps have a border of Barrier Peaks.

• Fire: Energy and Missile weapons can start forest fires. Moving or stopping in a forest fire adds a 6 points of heat per second to your mech, and the smoke adds three points on all hit rolls. Move at half normal rate.

• Cleared: Cannons and Missiles can clear trees from forest hexes. This is rough terrain, move at half normal rate. Cannons and Missiles can also be used to ’blow out’ forest fires.

• Burned: Once a forest fire has burned out the hex becomes rough terrain. Move at half normal rate.

Terrain plays an important part in combat strategy. A good position can mean the difference between winning and loosing. Use the terrain to your advantage. If you try to leave the map area you will be asked if your wish to ESCAPE from the combat. If you say Yes, your mech will be removed at the end of the move and you are out of the game. Forests are much harder to see through at ground level. If you are at ground level and the mech you are trying to see is also at ground level you can only see through three light woods or one light and one heavy woods.

12 Combat HEAT Affects

Almost everything you do in a Combat Mech creates heat. It is your worse enemy. Heat can slow down your mech, throw off your targeting computers, cause your mech to shutdown unexpectedly (that right, you’re a sitting duck), and even cause ammo to explode. Watch your heat. The Status display gives the amount of heat generated by each weapon in hundreds. Ten heat means it generates 1000 points of heat. The actual temperature is in the Status display. In addition, there is a small status light in the upper right corner of the weapons selection window. Green=OK, Blue=reduced movement/aim, Yellow=danger of shutdown, Red=danger of ammo exploding. Your Combat Mech will automatically shutdown if your heat goes over 3000. In addition, a damaged engine generates 5 points of heat per second for each hit. The only thing that protects a pilot from the heat is the life support system. If that is destroyed you are in big trouble.

13 Combat

There are two forms of combat, weapons combat and melee combat. Melee combat is hand to hand, punching, kicking and charging. Melee combat doesn’t generate any heat, but you may do damage to yourself and you have to be in the same hex or in the adjacent hex. When used to fight, your arms and legs have recycle times just like your other weapons. Legs take 25 seconds and arms take 10 seconds. Both legs must be ready in order to kick so you can only kick with one leg, however a kick does twice as much damage as a punch. If you miss a kick you have to make a Piloting Skill roll to keep from falling. On the other hand if you connect with a kick the kickee has to make a Piloting Skill roll. If you are jumping when you kick you will have to make a skill roll. If you fail your skill roll you will crash.

Damage is based on the weight of the attacking mech. A 50 ton mech can do 10 points of damage kicking and 5 points punching. To punch or kick simply select the arm or leg you want from the weapons selection window. The arm or leg must be ready and in working condition. If the status light to the right is RED it means that that limb is too damaged to be used in melee. The numbers after arm/leg status in the status display give an indication of that limbs condition. The first number is the modifier on ’to hit’ for weapons in the limb. The second number is a damage indicator. If it is over 6 the limb can’t be used. You can select multiple weapons so you could punch with both arms and kick with one leg in the same attack. You can not kick and fire any weapons in the same attack. You can not punch and fire any weapon located in that arm in the same attack.

When one mech attempts to move into the hex occupied by another mech, that is a charge. You must have both arms and both legs recycled and in good condition to walk or run charge. You must have both legs in good condition to jump charge. If your limbs aren’t ready or damaged you won’t be allowed to move into the other hex for the charge.

Charging can do massive damage. The damage is based on speed for walking and running charges. Your speed is the TIME it takes to walk or run one normal hex, divided into 100. This is modified by terrain and can be a fraction. The time is modified by heat and the damage to your legs. The damage is spread out in groups of 5 points. The charging mech takes damage equal to the weight of the mech charged divided by 10. A charge that hits requires both mechs to make a Skill roll at +2 or fall. You can NOT walk or run charge a mech that is down.

A jumping charge is called "Death from Above". Damage is the attacking mechs weight divided by 10, times three. An 80 ton mech jumping on you can do 24 points of damage. Damage is divided in half and each half applied to a random point in the upper torso, as though he kicked twice. The charging mech takes damage equal to his weight divided by 10, equally applied to his two legs.

To do a jumping charge, jump into the opponents hex and stop. All charges that hit require both mechs to make a Piloting Skill roll, the attacker at +2 and the defender at +2. Only the attacker makes a skill roll if the charge misses, but it is at +6. You CAN jump on a mech that is down.

If both mechs are jumping in the same hex there is a chance of a collision. This is not an attack and the computer rolls for whether a collision took place based on both pilots skill. If the mechs collide, both mechs must make a skill roll or fall and crash.

Melee is also dependant on the terrain. If you are too high above or too low below a mech you can’t melee. If you are one level above you can only kick him. If you are one level below you can only punch him.

Weapon combat is fairly selfexplanitory. Pick the weapons you want from the weapons selection window. That arms the weapons. You can select multiple weapons. When you select a weapon a pink number will appear beside the status indicator. This number is the order in which the weapons will be fired. Two seconds later the targeting computer will display a firing fan for that weapon. The firing fan will show your field of fire for chest mounted weapons and the weapons maximum range, and for some weapons the minimum range. Weapons mounted in the left arm can fire one hex side further to the left and weapons mounted in the right arm can fire one hex side further to the right. This means that arm mounted weapons can fire directly behind you if you swivel. THIS CAN BE VERY HANDY. You can fire inside the minimum range, but at a penalty to hit. The menu bar will list the last mech you fired at and several options. The left and right arrow scrolls through ALL enemy mechs. The "OK" option picks the current mechs as the target. The "Side" option allows you to select a mech on your side as the target. "Cancel" cancels the attack. "Hex" allows you to shoot at hexes.

The mech name might be colored. This is a status indicator for the mech. "Red" indicates that the mech is dead. "Blue" is destroyed. "Pink" is shutdown or the pilot is unconscious. "Yellow" is surrendered and "Green" means the mech has escaped off the screen. Select your target to activate the firing sequence. Two seconds later the weapons will fire. As each weapon comes up you will have the option to "Fire" the weapon, "Pass" on that one weapon, or "Modify" your gunnery skill roll. Hit locations will be flashed in the menu bar. If you miss a shot there is a chance that you could hit a mech next to your target by mistake.

Various objects are part of some scenarios. If your goal is to destroy the object you must be in the same hex with the object. You can then fire at it with weapons or kick and punch it. If your goal is to recover an object you will notice that all mechs with hands have an "H" in the status display next to the arm status. When you bring up the weapons menu those arms with hands will be listed as "Punch/Grab". If you select an arm with a hand the targeting menu will list the name of the object as an option. If you select the object your attack will give you a roll to pick up the object. If you are carrying an object your weapons menu will say "Punch/Drop". You can use the same procedure to drop an object with one or both hands. While carrying an object any stagger roll will also require a roll for dropping the object. Dropping an object will cause damage to the object.

Terrain is important in weapon combat. You can’t shoot through mountains so being on top of a mountain is a strategic advantage. Forests provide cover, but they can be set on fire or cleared of trees by concentrated fire. Water provide a good barrier from a charging mech. You can sometimes start fires by aiming at a forest hex. You can also blow out fires by shooting a fire hex with missiles or cannons. To shoot at a hex select HEX as the target of any weapon attack. You will then be asked to select the hex with the mouse. Click in the menu line to cancel the attack.

The main objective of combat is to eliminate the opponents armor so that your weapons can damage delicate internal components, like the pilot. Once the armor in a particular location is destroyed any further damage in that area hits something inside. The computer rolls a 2d6. If it is less than 8 the damage is applied to the Internal Structure of the mech. If it is 8 or higher one of the internal locations is hit. The internal locations contain such things as weapons, ammo, arm and leg joints, heat sinks, engines, gyros, PILOTS, etc. One or more hits will destroy an internal component.

The internal structure has points like the armor. These points are shown in the Status Display right after the area name and before the list of internal components. If the internal structure is reduced to zero that area is Destroyed and ALL components in it cease to function. If it is the head or limbs, the area is ripped off. If it is the center torso there is a 50/50 chance the engine will explode. Whether the engine explodes or not it stops working and the game is over for that mech. If a mechs engine explodes or its head is shot off the pilot will get a chance for an emergency eject. Don’t expect to survive, but you will have a chance.

14 A word about Orders

There are nine basic types of goals. These can be used in various combinations to create orders. At the end of the game your experience money and Combat Points will be determined on how well you completed the goal for your side. The criteria for each goal are:

• Kill Everything: You get 100% if all mechs are dead or disabled. You get 0% if any mechs survive or escaped.

• Scout/Patrol: At least one mech must survive/escape to answer ALL questions for the report at the end. If your side escapes leaving enemy mechs in control of the field your success is the ratio of the number of surviving mech divided into the number that went at LEAST half across the screen (to the dividing line).

• Recover, Find/Recover or Convoy Object: You fail if the object is destroyed or captured by the other side. If you get the object you a percentage based on how badly the object is damaged.

• Destroy or Find/Destroy Object: You get 100% if the object is destroyed.

• Convoy Across: Your success is measured by the number of mechs that successfully reach the opposite side.

• Hold Position: If at least half you all operational mechs have stayed in position (inside the circle) for the required time you are 100% successful (even if everyone dies after that).

15 Internal Components

• Weapons (W): Destroyed with any hit.

• Ammo (A): Destroyed with any hit. The ammo will blow up doing internal damage to the mech.

• Arm and Leg Actuators (x): Destroyed with any hit. Leg hits will require a Pilot Skill roll to keep from falling. The status of your arms and legs is given in the status display. The first number is the modifier to hit for weapons in the limb. The second number is a weighted actuator damage. If it is over 6 the limb is unusable.

• Engine (E): Each hit causes the engine to generate 5 points per second of extra heat. There is a chance on ANY hit that the fusion engine will blow up. The odds are very good by three hits and 100% by four hits.

• Gyro (G): A gyro hit adds 3 points to all Pilot Skill rolls.

• Jump Jets Ports (P): Destroyed with any hit. Each hit destroys that port and reduces your jump range one hex.

• Life Support (L): Any hit destroys the life support system for the mechs cockpit. The mech pilot will take one point of damage every 50 seconds that the heat is over 1500. If the heat is over 2500 he will take two points of damage.

• Sensors (S): The first hit on sensors add 2 points to all Gunnery Skill rolls to hit targets. A second hit destroys your sensors and you can not hit anything with your weapons. It is still possible to melee.

• Cockpit (C): Any hit in the cockpit kills the pilot. End of Game.

16 Hidden Commands

Hidden Commands There are several hidden commands in the "Setup" window. With the "Setup" window active

the following keys act as toggles:

"a" - Toggles Aimed Shots On/Off for ALL mechs. "r" - Toggles ALL Recycle Times to ZERO. "q" - Will Quit the game without saving mechs. This will leave you pilots "Out on Patrol".

They can be brought back with the HeadQuarters program, but it will cost you some money and will be counted as a loss on your combat record.

Concept Index 17

Concept Index

A

A word about Orders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 A word or two about memory . . . . . . . . . . . . . . . . . . . . . . . 4

C

Combat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Combat HEAT Affects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Copyright and Disclamier notice . . . . . . . . . . . . . . . . . . . . 1

H

Hidden Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

I

Internal Components . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

M

MechForce Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 MechForce Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Movement Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

P

Pulldown Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

S

Skill Rolls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

T

Terrain Types and Features . . . . . . . . . . . . . . . . . . . . . . . . 10 The Random Map Control Panel . . . . . . . . . . . . . . . . . . . . 9

W

Weapon Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 What is MechForce? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Table of Contents

1 Copyright and Disclamier notice . . . . . . . . . . . . . . . . . . . . 1 2 What is MechForce? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 3 MechForce Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4 A word or two about memory . . . . . . . . . . . . . . . . . . . . . . . 4 5 MechForce Documentation . . . . . . . . . . . . . . . . . . . . . . . . . . 4 6 Movement Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7 Weapon Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 8 Pulldown Menus. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 9 Skill Rolls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 10 The Random Map Control Panel . . . . . . . . . . . . . . . . . . 9 11 Terrain Types and Features . . . . . . . . . . . . . . . . . . . . . . . 10 12 Combat HEAT Affects . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 13 Combat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 14 A word about Orders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 15 Internal Components . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 16 Hidden Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 Concept Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17